# Diversity-Oriented Data Augmentation with Large Language Models

### v1

@@ Appendix D Augmentation Examples
Appendix E Parameter Sensitivity Study
@@
E.1 Number of Generated Sentences.
@@
E.2 Coreset Ratio.
We then use the MNLI dataset to study the effect of coreset ratio for data pruning and data augmentation. As presented in Figure 6b and Figure 6c, both data pruning and data augmentation favor a moderate ratio, where data pruning ratio is embodied by $r_{\text{prune}}/(r_{\text{prune}}+r_{\text{retain}}+r_{\text{augment}})$, and data augmentation is embodied by $r_{\text{augment}}/(r_{\text{retain}}+r_{\text{augment}})$.
E.3 Coreset Method.
The choice of coreset methods is another factor that influences the final performance. So, we also investigate how the performance changes when augmentation is applied to different coresets. Figure 7 shows that different datasets favor different coreset methods, however, in most cases, the suboptimal choices of coreset still outperform data augmentation without targeting coresets, and augmentation performance does not significantly degenerate on most suboptimal coresets. This result demonstrates our coreset-focused selective data augmentation method can benefit from appropriate coresets but is robust against suboptimal coresets.
@@
Figure 6: Parameter sensitivity study on generated sentence count, coreset ratio for data pruning and augmentation on MNLI dataset.
@@
Figure 8: Qwen performance on 12 datasets
Appendix F Full Results of LLM Architecture Exploration
@@
Table 7: Results on low-resource datasets (600), larger original subsets (1.2 K), and DoAug (600 Original + 600 Augmented Samples).
@@
(c) Distance
(d) Average accuracy
(e) Dispersion
@@
Appendix G Alleviating the Low-resource Problem

### v2

@@ Appendix E Augmentation Examples
Appendix F Diversity Evaluation
We present the original diversity scores of all methods in Table 7.
Distance Dispersion Radius Homogeneity Vocabulary 3-grams
Original 8.9440 0.1902 0.5354 0.8931 4734 23954
OCR 8.9514 0.1915 0.5340 0.8992 4793 25195
Keyboard 8.9474 0.1915 0.5334 0.8993 4829 25844
EDA 9.1324 0.2024 0.5287 0.9015 4974 29200
ADEA 8.9546 0.1926 0.5302 0.9073 4735 25596
BT 9.2111 0.2018 0.5347 0.8798 5193 30634
Unmask 9.0042 0.1932 0.5346 0.8994 4797 26774
AugGPT 9.1025 0.1956 0.5366 0.8757 5035 27420
Grammar 9.3905 0.2073 0.5373 0.8441 4895 30108
Spelling 9.0390 0.1949 0.5327 0.8996 5352 28154
Chain 9.1346 0.1955 0.5371 0.9076 5341 31543
Hint 9.3326 0.2043 0.5371 0.9016 5296 31649
Taboo 9.1241 0.1952 0.5368 0.9107 5172 30634
DoAug 9.6430 0.2180 0.5362 0.9095 5993 35308
Table 7: 6 diversity metrics averaged on 12 datasets.
Appendix G Measurements for Preference Dataset Construction and Diversity Sampling
In our final experimental settings, we use Euclidean distance to construct the preference dataset and sample the more diversified generations. An alternative to Euclidean distance is Cosine similarity. Before we settled on Euclidean distance, we examined and compared both Euclidean distance and Cosine similarity. We conduct the examination on a subset of 1000 samples from the original paraphrase dataset. We notice that when selecting the most diverse samples as “chosen samples”, in $97.1\%$ of the cases the two metrics yield the same samples. When selecting the most repetitive samples as “rejected samples”, in $97.8\%$ of the cases, the two metrics yield the same samples. For the total 5000 paraphrases (each sample contains 5 paraphrases), we also investigate the Pearson correlation between their Euclidean distance and dissimilarity (that is, 1 - Cosine similarity) compared with the original sentence, and find that the Pearson correlation is 0.96, indicating they are highly correlated. The relationship between Euclidean distance and dissimilarity for each paraphrase is shown in Figure 8a. Further, we observe that the sample distance is distributed more smoothly, suggesting samples of different diversity are more distinguishable, as shown in Figure 8b. Besides, Euclidean distance is more straightforward (higher distance is higher diversity) and easier to understand. Given the above arguments, we finally use Euclidean distance in our coding but expect the performance to be consistent if switching to Cosine similarity.
(a) Correlation
(b) Distribution
Figure 8: For (8a), we sample 500 points when plotting the diagram. For (8b), all scores are normalized to $[0,1]$.
Appendix H Full Results of Replacing DPO training
Detailed results of diversity in terms of 6 metrics are given in Figure 9.
(a) Distance
(b) Dispersion
(c) Radius
(d) Homogeneity
(e) Vocabulary
(f) 3-gram
Figure 9: Replacement study on DPO training.
Appendix I Parameter Sensitivity Study
@@
I.1 Number of Generated Sentences
@@
I.2 Coreset Ratio
We study the effect of the coreset ratio for data pruning and data augmentation to determine the best ratios. As presented in Figure 10b and Figure 10c, both data pruning and data augmentation favor a moderate ratio. The best performance occurs when the pruning ratio is $1/3$ and the augmentation ratio is $1/2$. In this way, $1/3$ of the total samples are pruned, and $2/3$ of the total samples are preserved. Then from the preserved $2/3$, we use $1/2$ of the remaining as seed samples for augmentation, which is $(2/3)\times(1/2)=1/3$ of the total samples. As a result, the final ratio $r_{\text{augment}}:r_{\text{retain}}:r_{\text{prune}}=1/3:1/3:1/3=1:1:1$.
@@ I.2 Coreset Ratio
Figure 10: Parameter sensitivity study on generated sentence count, coreset ratio for data pruning and augmentation on MNLI dataset. Data pruning ratio is embodied by $r_{\text{prune}}/(r_{\text{prune}}+r_{\text{retain}}+r_{\text{augment}})$, and data augmentation is embodied by $r_{\text{augment}}/(r_{\text{retain}}+r_{\text{augment}})$.
Appendix J Coreset Selection Methods
The choice of coreset methods is another factor that influences the final performance. So, we also investigate how the performance changes when augmentation is applied to different coresets, as presented in Figure 11. The result shows that different datasets favor different coreset methods, however, we notice that “variance” and “CCS w/ AUM” benefit most datasets (9 out of 12), and in most cases, the suboptimal choices of coresets still outperform data augmentation without targeting coresets, and augmentation performance does not significantly degenerate on most suboptimal coresets. This result demonstrates our coreset-focused selective data augmentation method can benefit from appropriate coresets but is robust against suboptimal coresets, and the “variance” and “CCS w/ AUM” can be used as the default coreset method.
@@
Appendix K Alleviating the Low-resource Problem
@@
Table 8: Results on low-resource datasets (800), larger original subsets (1.2 K), and DoAug (800 Original + 400 Augmented Samples).
Appendix L Full Results of LLM Architecture Exploration
@@
(c) Radius
(d) Homogeneity
(e) Vocabulary
(f) 3-gram
@@
Figure 13: Qwen performance on 12 datasets
(a) CSQA
(b) CODAH
(c) Multilingual
Figure 14: Performance improvement on two multi-choice-formed reasoning datasets and a multilingual text classification dataset.
Appendix M Applicability on More Datasets
To expand the evaluation scope and verify the broad application of DoAug, we additionally test the method on three datasets: CSQA, CODAH, and Multilingual, beyond the 12 English classification datasets in our main experiments. CSQA (CommonSenseQA) Talmor et al. (2019) and CODAH (COmmonsense Dataset Adversarially-authored by Humans) Chen et al. (2019) are two reasoning datasets. They are both in the form of multiple-choice questions. Multilingual Muennighoff et al. (2023); Barbieri et al. (2022); Enevoldsen et al. (2025) is a multilingual sentiment analysis dataset collected from Twitter. We filter the dataset to keep English, Spanish, German, French, Italian, and Portuguese samples, and remove Arabic samples because our augmenter is a Llama model and does not support Arabic. For the Multilingual dataset, we also include Code-Mixed, a data augmentation technique that switches some words to other languages and is designed especially for multilingual sentiment analysis tasks. As shown in Figure 14, DoAug also benefits multilingual datasets and reasoning tasks.